# Maximum Length RLL Sequences in de Bruijn Graph

### v1

@@ 2 The Maximum Length of an $(n,s)$-Sequence
In this section, we will present an upper bound on the length of $(n,s)$-sequences. Let ${\cal C}$ be a circuit of maximum length in $G_{n-1}$ which does not contain an edge with $s+1$ consecutive zeros in its representation. This circuit is a cycle in $G_{n}$ with no vertex with $s+1$ consecutive zeros in its representation and hence it represents an $(n,s)$-sequence. To find an upper bound on the length of ${\cal C}$ we will try to remove the edges from $G_{n-1}(s)$ which are not contained in ${\cal C}$. These edges are associated with vertices whose in-degree is not equal to their out-degree. The following lemmas are immediate consequences from the binary representation of the vertices, their in-edges, and out-edges.
@@
The vertices with in-degree one in $G_{n-1}(s)$ are all the vertices that start in their binary representation with exactly $s$ zeros followed by a one.
@@
The vertices with out-degree one in $G_{n-1}(s)$ are all the vertices that end in their binary representation with a one followed by exactly $s$ zeros.
@@
All the words in a necklace of order $n$ with at least two disjoint runs of $s+1$ or more zeros have at least one run with $s+1$ consecutive zeros.
@@
All the words of each necklace of order $n$ with at least two disjoint runs of $s+1$ zeros are not contained in the circuit ${\cal C}$.
@@
In $G_{n-1}$ a necklace of order $n$ that contains a unique run with more than $s$ zeros and the length of this run is $s+k$, $1\leq k\leq s$, contains $n-1-s+k$ edges with a run of more than $s$ zeros.
@@
Consider the edge $0^{k-1}X0^{s+1}$, $1\leq k\leq s$, where $X$ is an $s$-ones string of length $n-s-k$. This is the first edge in $G_{n-1}$ that has a run of $s+k$ zeros, associated with a necklace of order $n$, to be considered. The last edge with a run of more than $s$ zeros on this necklace to be considered is $0^{s+1}X0^{k-1}$. Now, in the path from the edge $0^{k-1}X0^{s+1}$ to the edge $0^{s+k}X$ in this necklace there are $1+k-1+n-s-k=n-s$ edges including the first edge and the last edge. All these edges have a run of zeros with a length greater than $s$. From the edge $0^{s+k}X$ to the edge $0^{s+1}X0^{k-1}$ in the necklace there are exactly $k-1$ edges including the last edge, excluding the first edge. Again, all these edges have a run of zeros with a length greater than $s$. Hence, in a necklace with a unique long run of more than $s$ zeros whose length is exactly $s+k$, the total number of edges with a run of at least $s+1$ zeros is $n-s+k-1$. ∎
@@
Any edge in $G_{n-1}(s)$ of the form $e=0^{s}X0^{k}$, where $1\leq k\leq s$, and $X$ is an $s$-ones string of length $n-s-k$, yields a path, which starts at the vertex $(0^{s}X0^{k-1})$, whose length is at least $s-k+1$, its edges do not have a run of more than $s$ zeros, and these edges are not contained in ${\cal C}$.
@@
For a vertex of the form $v=(0^{s}X0^{k-1})$, where $1\leq k\leq s$ and $X$ is a $s$-ones string, in $G_{n-1}(s)$, let ${\cal P}(v)$ denote the sub-path of length at least $s-k+1$ in $G_{n-1}(s)$, that starts in the vertex $v$ and ends at a vertex $v^{\prime}$ whose out-degree one (as described in the proof of Lemma 2.8) and whose edges are not contained in ${\cal C}$. The sub-path ${\cal P}(v)$ will be called the deleted path of $v$.
Corollary 2.9.
After all the deleted paths induced by the proof of Lemma 2.8 are removed from $G_{n-1}(s)$ the graph remains only with vertices whose in-degree equals to their out-degree.
Proof.
This is an immediate consequence of the proof of Lemma 2.8. It should be noted that for a vertex $v$ with two out-edges $v\rightarrow v_{1}$ and $v\rightarrow v_{2}$, one vertex (either $v_{1}$ or $v_{2}$) ends with a zero and hence the edge $e$ in Lemma 2.8 is taken without loss of generality as $e=0^{s}X0^{k}$, where $1\leq k\leq s$. ∎
Corollary 2.10.
There exists exactly one deleted path ${\cal P}(v)$ of length $s-k+1$ from the vertex $v=(0^{s}X0^{k-1})$, where $1\leq k\leq s$ and $X$ is a $s$-ones string.
@@
In $G_{n-1}(s)$, the first $s-i$ edges in a deleted path ${\cal P}(v)$ of $v=(0^{s}X_{1}0^{i})$ are disjoint from the first $s-j$ edges in a deleted path ${\cal P}(u)$ of $u=(0^{s}X_{2}0^{j})$, where $X_{1}$ and $X_{2}$ are two distinct $s$-ones strings and $i,j\geq 0$.
@@
The length of the circuit ${\cal C}$ (an $(n,s)$-sequence) in $G_{n-1}(s)$ is at most the number of edges in all the $(n,s)$-necklaces. If a circuit of such length exists it contains exactly all the words of the $(n,s)$-necklaces.
@@
Let $B$ be a necklace of order $n$ with a run of more than $s$ consecutive zeros. By Lemma 2.5 all the words in a necklace with a run of at least $12s+$ 1 zeros have at least one run with more than $s$ zeros and hence all the vertices of these necklaces are not contained in ${\cal C}$.
@@
If the longest run of consecutive zeros in $B$ is $s+k$, $1\leq k\leq s$, then by Lemma 2.7 there are $n-1-s+k$ edges in the necklace with a run of more than $s$ consecutive zeros. Such a necklace has an edge $e$ in $G_{n-1}(s)$ of the form $e=0^{s}X0^{k}$, where $1\leq k\leq s$ and $X$ is an $s$-ones string, whose first vertex is $v=(0^{s}X0^{k-1})$ and by Lemma 2.8 it yields a deleted path ${\cal P}(v)$ whose length is at least $s-k+1$. Together, $B$ ($n-1-s+k$ forbidden edges) and the associated deleted path ${\cal P}(v)$ (at least $s-k+1$ forbidden edges) have at least $(n-1-s+k)+(s-k+1)=n$ edges which are not contained in ${\cal C}$. By Lemma 2.11 all the deleted paths of different vertices are disjoint and hence the number of edges of $G_{n-1}(s)$ which are not contained in ${\cal C}$ is at least the number of edges in the necklaces of order $n$ which contain edges with a run of more than $s$ consecutive zeros.
Thus, the length of the circuit ${\cal C}$ in $G_{n-1}(s)$ is at most the number of edges in all the $(n,s)$-necklaces in $G_{n-1}(s)$.
For the fact that such a circuit cannot contain any $(n,s)$-word not from an $(n,s)$-necklace we have to add the consequence of Corollary 2.10. ∎
@@
What about the length of the maximum length trail ${\cal P}$ in $G_{n-1}(s)$? All the arguments we have used so far hold also for a trail with one exception. The trail can contain two vertices $v$ and $u$, where the in-degree of $v$ is one and its out-degree is two; the in-degree of $u$ is two and its out-degree is one. The first vertex in the trail is $v$ and the last vertex in the trail is $u$. We can apply the process in the proof of Lemma 2.8 on all the vertices in $G_{n-1}(s)$ except for $v$. After all the deleted paths are removed from $G_{n-1}(s)$ if we continue and apply the process in the proof of Lemma 2.8 on $v$ we will obtain a deleted path ${\cal P}(v)$ which starts in $v$ and ends at $u$. This deleted path can be added to the maximum length $(n,s)$-sequence to obtain a maximum length acyclic $(n,s)$-sequences. In other words, we have two vertices $v_{1}$ and $v_{2}$ for which we have the edges $v\rightarrow v_{1}$ and $v\rightarrow v_{2}$ in $G_{n-1}(s)$ and the in-degree of $v$ is one. Such vertices have the form $v=(0^{s}X)$, $v_{1}=(0^{s-1}X0)$, and $v_{2}=(0^{s-1}X1)$. Note, that $v$ is a vertex contained in the $(n,s)$-necklace $[10^{s}X]$ and $v$ is also a vertex contained in the necklace $[0^{s+1}X]$ which is not an $(n,s)$-necklace, but $s$ words of the necklace are $(n,s)$-words. The maximum length circuit in $G_{n-1}$ cannot contain both edges $v\rightarrow v_{1}$ and $v\rightarrow v_{2}$ since the in-degree of $v$ is one. But, a trail can contain both edges if one of them will be the first edge in the trail. In other words, if both edges are on the trail ${\cal P}$, then since they have the same unique predecessor in $G_{n-1}(s)$, it follows that one of them must be the first edge in ${\cal P}$. It can be only at the beginning of a sub-path of ${\cal P}$ which was excluded by Lemma 2.8. The length of this sub-path in Lemma 2.8 is $s$ and its structure is as follows:
@@
The length of a trail ${\cal P}$ of maximum length in $G_{n-1}(s)$ is at most the number of edges in all the $(n,s)$-necklaces plus $s$ (unless the $(n,s)$-necklaces contain all the $(n,s)$-words in which case the bound on the length of a trail is the same as the one on the length of a circuit).
@@
If $n=3$ and $s=2$, then an $(3,2)$-sequence of maximum length is
$0010111$
@@
If $n=5$ and $s=2$, then an $(5,2)$-sequence of maximum length has length 21 and one such sequence is.
001010011101011011111.001010011101011011111001010011101011011111~{}.001010011101011011111 .
An acyclic $(5,2)$-sequence has length 27 (4 additional bits for the acyclic representation and 2 additional bits from words which are not contained in $(5,2)$-necklaces) as follows
@@
A maximum length acyclic $(n,s)$-sequence (trail) in $G_{n-1}(s)$ can contain $s$ words which are not contained in a maximum length $(n,s)$-sequence. It was explained how to add them to a maximum length $(n,s)$-sequence. The maximum length trail can be constructed from an $(n,s)$-sequence in which one deleted path ${\cal P}(v)$ was taken from a necklace and it was added to obtain a maximum length $(n,s)$-sequence. It should be noted that the sequence starts and ends with the two possible shortest deleted paths. This is demonstrated in the following example.
@@
$00110101111101100101$
is a $(5,2)$-sequence of length 20 which contains two $(5,2)$-words $(00110)$ and $(01100)$ which are not contained in $(5,2)$-necklaces. This sequence falls short by one from the upper bound of a maximum length $(5,2)$-sequence. The deleted path which was removed from $G_{4}(2)$ is
@@
of length 27 is an acyclic $(5,2)$-sequence which contains 23 $(5,2)$-words, two of which are not contained in an $(5,2)$-necklace. This sequence attains the upper bound of a maximum length acyclic $(5,2)$-sequence. In this sequence the deleted path was added at the end of the sequence and hence all the words of the $(5,2)$-necklaces are contained in this trail.

### v2

@@ 2 The Maximum Length of an $(n,s)$-Sequence
In this section, we will present an upper bound on the length of $(n,s)$-sequences. To make the non-trivial proof simpler, it is broken to a sequence of claims which lead step-by-step to the main result. Let ${\cal C}$ be a circuit, of maximum length in $G_{n-1}$, which does not contain an edge with $s+1$ consecutive zeros in its representation. This circuit is a cycle in $G_{n}$ with no vertex having $s+1$ consecutive zeros in its representation and hence ${\cal C}$ represents an $(n,s)$-sequence. To find an upper bound on the length of ${\cal C}$ we will try to remove the edges from $G_{n-1}(s)$ which are not contained in ${\cal C}$. These edges are associated with vertices whose in-degree is not equal to their out-degree. The following lemmas are immediate consequences from the binary representation of the vertices, their in-edges, and out-edges.
@@
The vertices in $G_{n-1}(s)$ with in-degree one are those with the prefix $10^{s}$ 1.
@@
The vertices in $G_{n-1}(s)$ with out-degree one are those with the suffix $10^{s}$.
@@
All the words in a necklace of order $n$ with at least two disjoint runs of $s+1$ or more zeros have at least one acyclic run with $s+1$ consecutive zeros.
@@
All the words of each necklace of order $n$ with at least two disjoint runs of $s+1$ zeros are not contained in the circuit ${\cal C}$. In particular all the words of a degenerated necklace with at least one run of $s+1$ or more zeros are not contained in ${\cal C}$.
@@
In $G_{n-1}$ a necklace of order $n$ that contains a unique run with more than $s$ zeros and the length of this run is $s+k$, $1\leq k\leq s$, contains $n-s+k-1$ edges with a run of more than $s$ zeros.
@@
Consider a necklace of order $n$ that contains a unique run with $s+k$ zeros, $1\leq k\leq s$. The edges in the necklace that do not have a run with more than $s$ zeros are of the form $0^{i}X0^{s+k-i}$, $k\leq i\leq s$, where $X$ is an $s$-ones string of length $n-s-k$. There are $s-k+1$ such edges and a total of $n$ edges in the necklace and hence the total number of edges with a run of at least $s+1$ zeros is $n-(s-k+1)=n-s+k-1$. ∎
@@
Any edge in $G_{n-1}(s)$ of the form $e=0^{s}X0^{k}$, where $1\leq k\leq s$, and $X$ is an $s$-ones string of length $n-s-k$, yields a path, which starts at the vertex $v=(0^{s}X0^{k-1})$, whose length is at least $s-k+1$, its edges do not have a run of more than $s$ zeros, and these edges are not contained in ${\cal C}$.
@@
For a vertex of the form $v=(0^{s}X0^{k-1})$, where $1\leq k\leq s$ and $X$ is a $s$-ones string, in $G_{n-1}(s)$, let ${\cal P}(v)$ denote the sub-path of length $s-k+1$ of the path in $G_{n-1}(s)$, that starts in the vertex $v$ and ends at a vertex $v^{\prime}$ whose out-degree one (as described in the proof of Lemma 2.8) and whose edges are not contained in ${\cal C}$. The sub-path ${\cal P}(v)$ will be called the deleted path of $v$.
@@
Remark If $v=(0^{s}X0^{k-1})$, then its deleted path is of length $s-k+1$. We will prove later that this path contains exactly all the vertices that start with $v$ which were deleted as described in the proof of Lemma 2.8.
@@
In $G_{n-1}(s)$, the $s-i+1$ vertices in a deleted path ${\cal P}(v)$ of $v=(0^{s}X_{1}0^{i})$ are disjoint from the $s-j+1$ vertices in a deleted path ${\cal P}(u)$ of $u=(0^{s}X_{2}0^{j})$, where $X_{1}$ and $X_{2}$ are two distinct $s$-ones strings and $i,j\geq 0$.
@@
There exists unique deleted path ${\cal P}(v)$ of length $s-k+1$ from the vertex $v=(0^{s}X0^{k-1})$. All the deleted paths have distinct vertices.
@@
By the proof of Lemma 2.8 and the definition of the deleted path there exists a unique deleted path ${\cal P}(v)$. By Lemma 2.9 all these deleted paths have distinct vertices. ∎
The sequence of lemmas that were proved lead to the main results of this section.
Theorem 2.11.
For any $1\leq s<n$ we have that $m_{n,s}\leq\ell_{n,s}$.
Proof.
@@
Let ${\cal B}$ be a necklace of order $n$ with a run of more than $s$ consecutive zeros. By Lemma 2.5 all the words in a necklace with a run of at least $12s+$ 1 zeros have at least one run with more than $s$ zeros and hence all the vertices of these necklaces are not contained in ${\cal C}$. Similarly by Corollary 2.4 all the words of a degenerated necklace with at least one run of more than $s$ zeros have at least one run with more than $s$ zeros and hence all the vertices of these necklaces are not contained in ${\cal C}$.
@@
If the longest run of consecutive zeros in ${\cal B}$ is $s+k$, $1\leq k\leq s$, then by Lemma 2.7 there are $n-s+k-1$ edges in the necklace with a run of more than $s$ consecutive zeros. Such a necklace has an edge $e$ in $G_{n-1}(s)$ of the form $e=0^{s}X0^{k}$, where $1\leq k\leq s$ and $X$ is an $s$-ones string, whose first vertex is $v=(0^{s}X0^{k-1})$ and by Lemma 2.8 it yields a deleted path ${\cal P}(v)$ whose length is $s-k+1$. Together, ${\cal B}$ ($n-s+k-1$ forbidden edges) and the associated deleted path ${\cal P}(v)$ ($s-k+1$ forbidden edges) and possibly more edges that were deleted, we have at least $(n-s+k-1)+(s-k+1)=n$ edges which are not contained in ${\cal C}$. By Lemma 2.9 all these deleted paths have distinct vertices and hence the number of edges of $G_{n-1}(s)$ which are not contained in ${\cal C}$ is at least the number of edges in the necklaces of order $n$ which contain edges with a run of more than $s$ consecutive zeros.
Thus, the length of the circuit ${\cal C}$ in $G_{n-1}(s)$ is at most the number of edges in all the $(n,s)$-necklaces in $G_{n-1}(s)$, i.e., $m_{n,s}\leq\ell_{n,s}$. ∎
As will be described in Sections 3 and 5, it is well known that there exists $(n,s)$-sequence which contains all the words in the $(n,s)$-necklaces [17, 35] and hence we have the following consequence.
Corollary 2.12.
For any $1\leq s<n$ we have that $m_{n,s}=\ell_{n,s}$.
Corollary 2.13.
The deleted path ${\cal P}(v)$ of the vertex $v=(0^{s}X0^{k-1})$ is of length $s-k+1$. This path is the unique path from the vertex $v=(0^{s}X0^{k-1})$ to the vertex $(0^{k-1}X0^{s})$, where $1\leq k\leq s$ and $X$ is a $s$-ones string.
Corollary 2.14.
A circuit of length $\ell_{n,s}$ contains exactly all the words of the $(n,s)$-necklaces.
Proof.
For the analysis of Theorem 2.11 we have to add the consequences of Corollaries 2.10 and 2.13 (the uniqueness of the deleted paths) and observe that a circuit of length $\ell_{n,s}$ cannot contain any $(n,s)$-word which is not part of an $(n,s)$-necklace. ∎
@@
What about the length of the maximum length trail ${\cal P}$ in $G_{n-1}(s)$? All the arguments we have used so far hold also for a trail with one exception. The trail can contain two vertices $v$ and $u$, where the in-degree of $v$ is one and its out-degree is two; the in-degree of $u$ is two and its out-degree is one. In such a scenario, the first vertex in the trail is $v$ and the last vertex in the trail is $u$. In this trail we must have the two vertices $u$ and $v$ also as internal vertices (the trail starts at $v$ will arrive at $u$ continue to $v$ and ends at $u$). We can apply the process in the proof of Lemma 2.8 on all the vertices in $G_{n-1}(s)$ except for $v$. After all the deleted paths are removed from $G_{n-1}(s)$ if we continue and apply the process in the proof of Lemma 2.8 on $v$ we will obtain a deleted path ${\cal P}(v)$ which starts with $v$ and ends at $u$. This deleted path can be added to the maximum length cyclic $(n,s)$-sequence to obtain a maximum length acyclic $(n,s)$-sequences. It can be either at the beginning of the path or at the end of the path. In other words, for example we have two vertices $v_{1}$ and $v_{2}$ for which we have the edges $v\rightarrow v_{1}$ and $v\rightarrow v_{2}$ in $G_{n-1}(s)$ and the in-degree of $v$ is one. Such vertices have the form $v=(0^{s}X)$, $v_{1}=(0^{s-1}X0)$, and $v_{2}=(0^{s-1}X1)$. Note, that $v$ (a word of length $n-1$) is a vertex contained in the $(n,s)$-necklace $[10^{s}X]$ and $v$ is also a vertex contained in the necklace $[0^{s+1}X]$ which is not an $(n,s)$-necklace, but $s$ words of the necklace are $(n,s)$-words. The maximum length circuit in $G_{n-1}(s)$ cannot contain both edges $v\rightarrow v_{1}$ and $v\rightarrow v_{2}$ since the in-degree of $v$ is one. But, a trail can contain both edges if one of them will be the first edge in the trail. In other words, if both edges are on the trail ${\cal P}$, then since they have the same unique predecessor in $G_{n-1}(s)$, it follows that one of them must be the first edge in ${\cal P}$. It can be only at the beginning of a sub-path of ${\cal P}$ which was excluded by Lemma 2.8. The length of this sub-path in Lemma 2.8 is $s$ and its structure is as follows:
@@
If $s<n-1$, then the length of a trail ${\cal P}$ of maximum length in $G_{n-1}(s)$ is at most $\ell_{n,s}+s$ (if $s\geq n-1$, then its length is is at most $\ell_{n,s}$). The length of the associated acyclic $(n,s)$-sequence is $\ell_{n,s}+s+n-1$. In any such sequence we have all the words in the $(n,s)$-necklaces and one deleted path of length $s$.
@@
If $n=3$ and $s=2$, then a $(3,2)$-sequence of maximum length is
$[0010111]$
@@
If $n=5$ and $s=2$, then an $(5,2)$-sequence of maximum length has length 21 and one such sequence is
$[110010100111010110111]~{}.$
An acyclic $(5,2)$-sequence has length 27 (4 additional bits, $1100$ which are also the first 4 bits for the acyclic representation, and 2 additional bits from 2 words, $(00110)$ and $(01100)$, which are not contained in $(5,2)$-necklaces) as follows
@@
A maximum length acyclic $(n,s)$-sequence (trail) in $G_{n-1}(s)$ contains $s$ words which are not contained in a maximum length $(n,s)$-sequence (circuit). It was explained how to add them to a maximum length $(n,s)$-sequence. The maximum length trail can be constructed from a maximum $(n,s)$-sequence. One deleted path ${\cal P}(v)$ of length $s$ is taken from a necklace and it is added to obtain a maximum length acyclic $(n,s)$-sequence. It should be noted that the maximum length $(n,s)$-sequence either starts or ends with a deleted path. If words of a deleted path are added in other places a large cycle can be obtain, but it will fall short of $\ell_{n,s}$. This is demonstrated in the following example.
@@
$[00110101111101100101]$
of Example 4 is a $(5,2)$-sequence of length 20 which contains two $(5,2)$-words $(00110)$ and $(01100)$ which are not contained in $(5,2)$-necklaces. This sequence falls short by one from the upper bound of a maximum length $(5,2)$-sequence. The deleted path which was removed from $G_{4}(2)$ is
@@
of length 27 is an acyclic $(5,2)$-sequence which contains 23 $(5,2)$-words, two of which are not contained in an $(5,2)$-necklace. This sequence attains the upper bound of a maximum length acyclic $(5,2)$-sequence presented in Theorem 2.15. In this sequence the deleted path was added at the end of the sequence and hence all the words of the $(5,2)$-necklaces are contained in this trail.
@@ 3 Construction of Sequences of Maximum Length
In this section, we will concentrate on constructing $(n,s)$-sequences of maximum length. We will show that there exist many $(n,s)$-sequences which attain the upper bound of Theorem 2.11 and hence these sequences are maximum length $(n,s)$-sequences. In the literature there are many efficient algorithm to generate one de Bruijn sequence, but these algorithms cannot be used to generate many de Bruijn sequences with the same efficiency.